OpenStudy (anonymous):

Write the slope-intercept form of the line that passes through the point (1, 0) and is parallel to x - y = 7.

OpenStudy (studygurl14):

First step: Put x - y = 7 into slope-intercept form.

OpenStudy (ribhu):

line is parallel to the line x-y = 7 so slope of the line is 1

OpenStudy (ribhu):

now the slope intercept equation can be written as (y - 0) = slope (=1) (x-1) y=x-1

OpenStudy (anonymous):

So it's y=x-1?